A laser used to weld detached retinas emitslight with a wavelength 652 nm in pulses thatare of 20ms duration. The average powerduring each pulse is 0.6W. The energy in eachpulse and in a single photon areAns 7.5*10^16 ev , 1.9 ev

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
E(ev) = 1240/652 = 1.9 ev (in a single photon)
Energy in 20 ms = 20 * 0.6 * 10^-3 = 1.2 * 10^-2 joules
= 1.2 * 10^-2/ 1.6 * 10^-19
= 7.5 * 10^16 ev